Convert the following integral to an integral in polar coordinates. Je V x2 + y2 d xd y where Ris 45x+yºs -) SºS *rar de
Answer #1

DE Y cose y = rsimo dady = Jacotian oz (ou FYSIDE yasing = Jae dady = ( drdo I coso sino 1 -Yine ruose = rose tosiño = r unotdady= Jouy). drdo = y dede ( 1218 andy = rando 21 I r drdo ET (81-16 do ut com cro) do = + [50-s de ses (0) = 0 (211-07 = 6